Supernormal Growth [LO1]

Marcel Co. is growing quickly. Dividends are expected to grow at a 24 percent rate for the next 3 years, with the growth rate reducing to only a constant 5 percent thereafter.

Required:

If the required return is 13 percent and the company just paid a $2.50 dividend, what is the current share price? Note: since the dividend at time 0 of $2.50 has just been paid, do not include it in the price at time 0.  (Do not round your intermediate calculations.)

Multiple Choice

  • $52.42

  • $47.43

  • $51.37

  • $53.46

  • $49.11
Answer #1

D1=(2.5*1.24)=3.1

D2=(3.1*1.24)=3.844

D3=(3.844*1.24)=4.76656

Value after year 3=(D3*Growth rate)/(Required rate-Growth rate)

=(4.76656*1.05)/(0.13-0.05)

=62.5611

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=3.1/1.13+3.844/1.13^2+4.76656/1.13^3+62.5611/1.13^3

=$52.42(Approx).